Grade 11Algebra

If three non-zero distinct real numbers form an arithmetic progression and the squares of these numbers taken in the same order constitute a geometric progression, then find the sum of all possible common ratios of the geometric progression.

Jonathan Jackson
11 Years agoGrade 11

1 Answer

Harsh Patodia
ApprovedApproved Tutor Answer11 Years ago
Hi Student

Let the numbers be a-d,a,a+d
Squares of these number form a GP
(a-d)2 , a2 , (a+d)2

=> (a2)2 = (a-d)2(a+d)2
= > a4 = ( a2-d2)2
=> a2 = a2-d2 or a2= d2-a2
=> d=0 or d= ± \sqrt{2}a
From 1st case not possible as the numbers are distinct

r= a2./ (a-d)2 also r = (a+d)2/a2
substitue d as obtained above and get the possible values of r .